# Explainable Multi-Modal Data Exploration in Natural Language via LLM Agent

### v1

@@ 4.3. Results on the EHRXQA Dataset
Table 2. Performance metrics of NeuralSQL (zero-shot and few-shot) and XMODE (zero-shot) on the EHRXQA dataset.
(a) CAESURA
(b) XMODE
@@
Figure 7. Error analysis of CAESURA (urban2023caesura, ) (a) and XMODE (b) on the ArtWork dataset across different steps.
Table 2 demonstrates the experimental results of XMODE against NeuralSQL on the EHRXQA dataset. Our evaluation encompasses three scope categories: single-table queries with one image (Image Single-1), single-table queries with two images (Image Single-2), and multiple-table queries with single images (Image+Table Single). XMODE demonstrates robust performance across all evaluation metrics, achieving an overall accuracy of 51.00%. Notably, XMODE excels in handling multiple-table scenarios, where it achieves 77.50% accuracy, significantly outperforming NeuralSQL’s 47.50% in the 10-shot setting. For single-table queries, XMODE shows strong performance with 43.33% accuracy on two-image queries, though it achieves a slightly lower score (23.33%) compared to NeuralSQL’s 10-shot performance (26.67%) on single-image queries.
When examining the output types, XMODE exhibits particularly strong performance on binary questions, achieving 74.00% accuracy compared to NeuralSQL’s 48.00%. For categorical questions, both systems show lower performance, with XMODE reaching 28.00% and NeuralSQL achieving 18.00% in the 10-shot setting.
A key distinguishing feature of XMODE is its comprehensive functionality beyond raw accuracy. Unlike NeuralSQL, XMODE generates executable plans with 98% coverage, provides explanations for traceability of final outputs, and supports dynamic replanning capabilities. In contrast, NeuralSQL, even in its 10-shot configuration, lacks these additional features and shows no performance in the zero-shot setting across all metrics. These results highlight XMODE’s effectiveness as a more complete solution for EHRXQA tasks, particularly in complex scenarios involving multiple tables and binary decisions, while also offering important auxiliary features for practical deployment.
4.4. Error Analysis
In this section, we conduct a comprehensive analysis of errors encountered during the evaluation process. These errors are systematically classified into the following categories:
  • Planning Errors: These errors stem from incorrect or incomplete task planning, such as task decomposition, the generation of completely faulty natural language questions, etc.
  • Text-to-SQL Errors: Errors where the generated SQL fails to accurately retrieve the intended data.
  • Image Analysis Inaccuracy: Errors caused by inaccurate outputs from the image analysis model, even when the underlying task plan is correct.
  • Plot Generation Errors: Errors where plots are completely not generated, partially generated or incorrectly visualized, thereby failing to meet expected outcomes.
To systematically analyze key issues, we prioritize the identified categories based on their inter-dependencies during task execution. The priority sequence of these categories is defined as follows: task planning ¿ text-to-SQL generation ¿ Image analysis ¿ plot generation. Only the first affected category is considered if an error occurs at any stage, which may involve issues across multiple categories. For instance, if an error is detected during the planning phase but the subsequent tasks are successful, and another error occurs at the later plot generation stage, only the error in the planning phase is counted. In this case, the sample is classified under the Planning Error category.
This approach to error analysis is grounded in the logical dependency structure of the tasks. Since each task is a prerequisite for the succeeding one, a failure in an earlier task renders the success of subsequent tasks irrelevant to the overall reasoning process. As a result, errors are attributed to the earliest point of failure better to reflect the hierarchical nature of the task dependencies, thereby facilitating targeted optimization.
4.4.1. Error Analysis on the Artwork Dataset.
@@ 4.4. Error Analysis
For 11 errors in multi-modal questions, five are related to single-value outputs, four to plots, and three to data structures. All of these errors are attributed to incorrect outputs generated by the image analysis model. After further research, we found two ambiguous tasks in classifying the error categories. (1) Plot the number of paintings that depict war for each year and (2) What is depicted on the oldest religious artwork in the database? Both tasks failed due to improperly parsed sub question for the image analysis task, specifically the oversimplified term “war.” While this term is semantically related to the correct natural language question, “Does the image depict war?”, it does not fully capture the intent of the task. As a result, it cannot be classified as a completely faulty question. Notably, the XMODE model generated correct results for these tasks, underscoring the limitations of CAESURA’s approach in handling subtle semantic distinctions.
@@
In questions which require a parallel workflow - including two data structure , two plot—plot, and two plot—data structure outputs — errors are observed at the early planning stage. Our analysis reveals that CAESURA encounters significant challenges in generating accurate plans for embarrassingly parallel tasks. For two of these tasks, the system fails to generate any plan at all. For the remaining four tasks, CAESURA can provide partial results for some subtasks, but other subtasks are left unanswered, reflecting a broader issue in its ability to manage parallel planning. Our XMODE system successfully generates the appropriate plans for all tasks. In addition, all text-to-SQL steps , data preparation pipelines, and plot outputs, where required, are validated as correct. As illustrated in Figure 7(b), the only source of errors is the inaccurate output of the image analysis model, which accounted for 11 errors.No other errors are located in the text-to-SQL task, plot generation, or task planning deficiencies. This analysis highlights the image analysis model as the bottleneck in system performance, underscoring the need for further refinement in its predictive accuracy.
4.4.2. Error Analysis on the EHRXQA Dataset
@@
Since NeuralSQL is a one-step approach lacking task planning and explainability, we are unable to localize the source of errors as systematically as in the XMODE or CAESURA systems. Consequently, we focus our error analysis solely on the XMODE system using the EHRXQA dataset.
@@
Figure 8 presents the distribution of 49 errors across various steps, categorized by their respective scopes: Image Single-1 (23 errors), Image Single-2 (17 errors), and Image+Table Single (9 errors). Among these, 36 errors are associated with the categorical scope, with 20 attributed to Image Single-1 and 16 to Image Single-2. In contrast, errors linked to the binary output type are primarily found in the Image+Table Single scope. Specifically, Image Single-1 contributes three binary errors, Image Single-2 accounts for one, and Image+Table Single includes nine, summing up to 13 binary errors out of the total 49. Considering the uneven distribution of errors across various output types and scopes, we identified inaccurate image analysis — primarily driven by the M3AE model (10.1007/978-3-031-16443-9_65, ) — as the main source of errors. Our analysis reveals that errors linked to categorical output types (36) are nearly three times higher than those associated with binary output types (13). This suggests that the error pattern is less related to the task difficulty across different scopes and more influenced by the output type, as binary questions demonstrate a statistically higher success rate compared to categorical ones. Notably, the Image + Table Single scope exclusively utilizes binary output types.
@@
To gain a deeper understanding, a step-by-step error analysis reveals that out of the 23 errors in the Image Single-1 scope, 22 are due to inaccuracies in image analysis, while only one is related to a misstep in the text-to-SQL process. The specific question text for this case is: “Catalog all the anatomical findings seen in the image, given the first study of patient 11801290 on the first hospital visit.” The generated SQL query fails to include the condition specifying the first study, resulting in an incorrect output. In the Image Single-2 category, 16 out of 17 total errors are due to inaccurate image analysis, with one error attributed to the text-to-SQL step. The specific query in question is: “Does the second-to-last study of patient 16345504 this year reveal still-present fluid overload/heart failure in the right lung compared to the first study this year?”. The text-to-SQL task fails to correctly retrieve the first and last study of this year as required, instead erroneously returning multiple studies from the current year. In the Image+Table Single scope, all nine errors involve binary output types. Of these, six result from inaccurate image analysis, one from incomplete planning, and two from an incorrect text-to-SQL step. The error caused by incomplete planning occurs with the question: “Did patient 19055351 undergo the combined right and left heart cardiac catheterization procedure within the same month after a chest x-ray revealed any anatomical findings until 2104?”. In this case, the plan omits the necessary image analysis step, leading to an incorrect final output. During the reasoning stage, instances were identified where an empty output produced a no response that coincidentally aligned with the ground truth. However, XMODE’s explainability highlights this as a misclassification, as the absence of output was not due to correct reasoning.
@@
Two errors in the Image+Table Single category are attributed to text-to-SQL misbehavior. The specific questions causing these errors are: ”Was patient 12724975 diagnosed with hypoxemia until 1 year ago, and did a chest x-ray reveal any tubes/lines in the abdomen during the same period?” and ”Was patient 10762986 diagnosed with a personal history of tobacco use within the same month after a chest x-ray showing any abnormalities in the aortic arch until 1 year ago?” In both cases, the SQL queries fail to correctly apply the condition (since current time) until 1 year ago, instead treating 1 year ago as a fixed point in time.
@@
Figure 8. Error analysis of XMODE on the EHRXQA (bae2024ehrxqa, ) dataset across different steps.
5. Conclusions
We demonstrated that multi-agent collaboration using large language models, such as GPT-4, offers a promising approach for explainable multi-modal data exploration in natural language. Our experimental evaluation against two state-of-the-art systems on two different datasets with tabular and image data shows that XMODE not only performs the task of multi-modal data exploration with higher accuracy but also faster due to smart re-planning and parallel execution. Moreover, XMODE also provides detailed explanations and reasoning which makes it transparent and supports the end user to better understand and verify the results. The main findings from our experiments are that the text-to-SQL task shows high accuracy, while the image analysis task only shows limited accuracy. Hence, future work on multi-modal data exploration should focus on improving the accuracy of the image interpretation and understanding models. Potential avenues for research are better alignment approaches between tabular and image data or iterative prompt engineering with natural language question re-writing to better probe the image search space. Another promising approach is to focus on human in the loop approaches where the system and the humans solve tasks jointly.
6. Acknowledgements
We want to thank Magda Balazinska from the University of Washington for fruitful discussions on the initial idea of the XMODE system. We also want to thank the CAESURA (urban2023caesura, ) and EHRXQA (bae2024ehrxqa, ) team for providing their expertise in running their respective systems.

### v2

@@ 4.4 Results on the Benchmark Datasets
Table 2: Performance metrics of NeuralSQL (zero-shot and few-shot) and M2EX (zero-shot) on EHRXQA. Planner coverage (Generated Plan): 98%.
Dataset System Tasks Errors Dominant Error Source
ArtWork CAESURA 30 20 Faulty plans; VQA errors
M2EX 30 11 VQA errors only
RotoWire CAESURA 12 9 Text analysis; SQL faults
M2EX 12 4 Text analysis
EHRXQA NeuralSQL 100 67 N/A – no plan output
M2EX 100 49 VQA errors only
Table 3: Top-level error breakdown. See App. E for details.
Results on the ArtWork and RotoWire Datasets
Table 1 shows M2EX outperforms CAESURA by 30% on the ArtWork and by ca. 42% on the RotoWire datasets in accuracy, with advantages in both single- and multi-modality queries. Efficiency-wise, M2EX excels on ArtWork with fewer steps, lower latency, and reduced costs. On RotoWire, despite higher token usage and costs due to advanced text analysis, M2EX maintains superior accuracy. Additionally, M2EX supports re-planning and offers better explanations, features absent in CAESURA.
Results on the EHRXQA Dataset In Table 3, M2EX outperforms NeuralSQL in overall accuracy (51.00% vs. 33.00% in 10-shot) on the EHRXQA dataset, especially in multiple-table queries (77.50% vs. 47.50%) and binary questions (74.00% vs. 48.00%). Additionally, M2EX provides plan generation (98% coverage), explanations, and replanning—features that NeuralSQL lacks. Metrics like steps, tokens, and latency are excluded since NeuralSQL generates answers directly without intermediate steps, unlike M2EX’s transparent workflow. We exclude CAESURA from the EHRXQA experiments due to its inefficiency with EHRXQA’s complex schema. While CAESURA is intended to be a general-purpose multi-modal system, it processes the relational database through multiple steps, examining each table and relationship sequentially. This limitation introduces significant overhead when handling the complex data schema of the EHRXQA dataset (there are 18 tables) during its discovery phase. Consequently, reproducing CAESURA on EHRXQA questions fails to perform inferences at the early stages of the planning phase, ultimately terminating after exceeding the maximum number of allowed attempts.
4.5 Error Analysis
We evaluate system errors across three datasets: ArtWork, RotoWire, and EHRXQA, identifying key bottlenecks and component failures (see Table 3 and detailed breakdown in Appendix E, Fig. 6). On the ArtWork dataset, CAESURA exhibits 20 errors out of 30 tasks, mainly due to faulty planning in sequential workflows and incorrect outputs from the image analysis module. Multi-modal tasks involving plot and data structure outputs are particularly error-prone, especially in parallel workflows where planning failures are common. By contrast, M2EX achieves full planning success, with image interpretation errors being the only significant issue. In the RotoWire dataset, CAESURA fails on 9 of 12 tasks due to text analysis failures and SQL generation flaws. M2EX resolves all single-modal tasks but faces 4 errors in multi-modal tasks, again tied to text interpretation. These patterns highlight M2EX’s robustness in planning and execution while exposing shared weaknesses in text and image understanding across systems.
For the EHRXQA dataset, we focus solely on M2EX due to NeuralSQL’s lack of interpretable planning. Of 49 errors, 36 arise in categorical tasks, indicating a strong link between output type and model performance. Most failures originate from inaccurate image analysis by the M3AE model. These results emphasize the need for improved image understanding, especially for categorical reasoning, alongside stronger planning and SQL components. See Appendix E for full error analysis.
5 Conclusions
In this paper, we show that multi-agent collaboration via LLMs (GPT-4o) offers a powerful approach to multi-modal data exploration in natural language. Our system, M2EX, outperforms prior methods across datasets with tabular, text, and image data by leveraging smart re-planning, parallel execution, and transparent, explainable workflows. It blends accuracy, efficiency, and user-centric design, marking a significant advance in multi-modal data exploration, with strong performance in text-to-SQL tasks and potential for further enhancement in image reasoning and workflow optimization.
Future work will focus on better data alignment, prompt design, planning efficiency, and scaling to larger datasets and new modalities such as video and human-in-the-loop interaction.
Limitations
Despite M2EX’s overall superior performance, several limitations remain. Most notably, the system’s reliance on image analysis introduces a consistent source of error, particularly in tasks involving categorical outputs. The M3AE model often fails to capture subtle visual distinctions, which disproportionately affects the accuracy of multi-modal tasks. We did not explore alternative image processing approaches, as improving the visual pipeline was not the primary objective of this study. Instead, we adopted visual models commonly used in prior work to ensure a fair and consistent basis for comparison. Similarly, we restricted our language model experiments to GPT-4o to both showcase our proposed methods and maintain comparability with recent studies.
Additionally, although M2EX successfully generates plans for all tasks, its performance still hinges on accurate text interpretation. In the RotoWire dataset, for example, errors in multi-modal questions were largely driven by flawed text comprehension, revealing a vulnerability in the language understanding pipeline.
Finally, the system exhibits a performance gap between binary and categorical tasks, suggesting that output type complexity influences success rates. These findings indicate that further improvements are needed in visual reasoning, nuanced language understanding, and output-type generalization.
Acknowledgments
This project has received funding from the Hasler Foundation Project ID: 2024-05-21-76, DIZH Project-Call 2024.1, ID:9, and the DataGEMS, funded by European Union’s Horizon Europe Research and Innovation programme, under grant agreement No 101188416.
Appendix A M2EX on ArtWork
Figure 3: M2EX framework on ArtWork  (Urban and Binnig, 2024) with an example of processing a multi-modal query. The query is automatically decomposed into various components such as text2SQL, and image analysis which can be inspected by the user for explainability.
Appendix B Prompts
Planner Prompt / Replanning Prompt [SYSTEM]: Given a user question and a database schema, analyze the question to identify and break it down into relevant sub-questions. Determine which tools (e.g., {tool_names}) are appropriate for answering each sub-question based on the available database information and tools. Decompose the user question into sub-questions that capture all elements of the question’s intent. This includes identifying the main objective, relevant sub-questions, necessary background information, assumptions, and any secondary requirements. Ensure that no part of the original question’s intent is omitted, and create a list of individual steps to answer the question fully and accurately using tools. You may need to use one tool multiple times to answer the original question. First, you should begin by thoroughly analyzing the user’s main question. It’s important to understand the key components and objectives within the query. Next, you must review the provided database schema. This involves examining the tables, fields, and relationships within the database to identify which parts of the schema are relevant to the user’s question and contribute to a set of sub-questions. For each sub-question, provide all the required information that may required in other tasks. In order to find this information look at the user question and the database information. Each sub-question or step should focus exclusively on a single task. Each sub-question should be a textual question. Don’t generate a code as a sub-question. Create a plan to solve it with the utmost parallelizability. Each plan should comprise an action from the following {num_tools} types: {tool_descriptions} {num_tools}. join(): Collects and combines results from prior actions. - An LLM agent is called upon invoking join() to either finalize the user query or wait until the plans are executed. - join should always be the last action in the plan, and will be called in two scenarios: (a) if the answer can be determined by gathering the outputs from tasks to generate the final response. (b) if the answer cannot be determined in the planning phase before you execute the plans. Guidelines: - Each action described above contains input/output types and descriptions. - You must strictly adhere to the input and output types for each action. - The action descriptions contain the guidelines. You MUST strictly follow those guidelines when you use the actions. - Each action in the plan should strictly be one of the above types. Follow the Python conventions for each action. - Each action MUST have a unique ID, which is strictly increasing. - Inputs for actions can either be constants or outputs from preceding actions. In the latter case, use the format \$id to denote the ID of the previous action whose output will be the input. - If there is an input from preceding actions, always point its id as ‘\$id‘ in the context of the action - Always call join as the last action in the plan. Say ’<END_OF_PLAN>’ after you call join. - Ensure the plan maximizes parallelizability. - Only use the provided action types. If a query cannot be addressed using these, invoke the join action for the next steps. - Never introduce new actions other than the ones provided. {list of usecase-specific business rules} [USER]:{state} [SYSTEM]: Remember, ONLY respond with the task list in the correct format! E.g.: idx. tool(arg_name=args),
Prompt for Decision Making [SYSTEM]: Solve a question answering task. Here are some guidelines: - In the Assistant Scratchpad, you will be given results of a plan you have executed to answer the user’s question. - Thought needs to reason about the question based on the Observations in 1-2 sentences. - Ignore irrelevant action results. - If the required information is present, give a concise but complete and helpful answer to the user’s question. - If you are unable to give a satisfactory finishing answer, replan to get the required information. Respond in the following format: Thought: <reason about the task results and whether you have sufficient information to answer the question> Action: <action to take> - If an error occurs during previous actions, replan and take corrective measures to obtain the required information. - Ensure that you consider errors in all the previous steps, and try to replan accordingly. - Ensure the final answer is provided in a structured format as JSON as follows: {{’Summary’: <concise summary of the answer>, ’details’: <detailed explanation and supporting information>, ’source’: <source of the information or how it was obtained>, ’inference’:<your final inference as YES, No, or list of requested information without any extra information which you can take from the ‘labels‘ as given below>, ’extra explanation’:<put here the extra information that you don’t provide in inference >, }} In the ‘inference‘ do not provide additional explanation or description. Put them in ‘extra explanation‘. Available actions: (1) Finish (the final answer to return to the user): returns the answer and finishes the task. (2) Replan(the reasoning and other information that will help you plan again. Can be a line of any length): instructs why we must replan. [USER]: {state} [SYSTEM]: Using the above previous actions, decide whether to replan or finish. If all the required information is present, you may finish. Consider replanning for data_preparation task if you want to structure the response in a proper way. If you have made many attempts to find the information without success, admit so and respond with whatever information you have gathered so the user can work well with you. Do not generate a response based on the sample data (assumption). If you failed after multiple attempts, you can finish and explain the reason.
Prompt for text2SQL [SYSTEM]: You are a database expert. Generate a SQL query given the following user question, database information and other context that you receive. You should analyse the question, context and database schema and come up with the executable sqlite3 query. Provide all the required information in the SQL code to answer the original user question that may required in other tasks utilizing the relevant database schema. Ensure you include all necessary information, including columns used for filtering, especially when the task involves plotting or data exploration. This must be taken into account when performing any time-based data queries or analyses. Translate a text question into a SQL query that can be executed on the SQLite database. You should stick to the available schema including tables and columns in the database and should not bring any new tables or columns. [USER]: {text2SQL task description}, {db schema}
Prompt for text_analysis [SYSTEM]: You are a text analysis assistant. Analyze the provided question and report to answer the question. Only answer the question and don’t provide extra information in your answer. In your answer, be concrete and use None if you can’t find the answer in the report. The output should be in the format: {{’reasoning’: ’...’, ’answer’: ’...’}} [USER]: {text analysis task description}, {text}
Prompt for data_preparation [SYSTEM]: You are a data preparation and processing assistant. Create a proper structure for the provided data from the previous steps to answer the request. - If the required information has not found in the provided data, ask for replanning and ask from previous tools to include the missing information. - You should include all the input data in the code, and prevent of ignoring them by ‘# ... (rest of the data)‘. - You should provide a name or caption for each value in the final output considering the question and the input context." - Don’t create any sample data in order to answer to the user question. - You should print the final data structure. - You should save the final data structure at the specified path with a proper filename. - You should output the final data structure as a final output. [USER]: {data preparation task description}, {result from previous task}
Prompt for data_plotting [SYSTEM]: You are a data plotting assistant. Plot the provided data from the previous steps to answer the question. - Analyze the user’s request and input data to determine the most suitable type of visualization/plot that also can be understood by the simple user. - If the required information has not been found in the provided data, ask for replanning and ask from previous tools to include the missing information. - Don’t create any sample data in order to answer to the user question. - You should save the generated plot at the specified path with the proper filename and .png extension. [USER]: {data plotting task description}, {data}
Appendix C Tools, Models, and Prompts by Subtask
Task Tool / Model Prompt Type
Text-to-SQL translation GPT-4o text2SQL prompt
Text analysis GPT-4o text_analysis prompt
ArtWork VQA BLIP-2 no prompt
Medical image (EHRXQA) VQA M3AE no prompt
Data preparation GPT-4o and Python (Pandas) data_preparation pormpt and Code via LLM output
Plot generation GPT-4o and Matplotlib + Pandas data_plotting prompt and Chart Code via LLM output
DAG construction (planning/replannig) GPT-4o (Planner loop) planner Prompt / replanning prompt
Decision Making GPT-4o decision making prompt
Table 4: Subtasks, their associated tools/models, and prompt styles used in M2EX. Most tool invocations are zero-shot or template-based.
Appendix D Optimizations in M2EX Explained with Examples
To better demonstrate advantages of M2EX, we provide several examples (see Figures 3 and 4) across three key aspects: explanations, smart replanning, and parallel planning. The following examples provide a detailed illustration of these three aspects.
3) Decision Making: When the tasks are completed, the model reflects on its work and provides a final output based on its thought as Summary:"The number of paintings depicting war has been plotted for the 16th and 18th centuries.", "Details": "The analysis identified 1 painting from the 16th century and 2 paintings from the 18th century that depict war. The plot visualizes these findings. [..]". Throughout the workflow, the model demonstrates a commitment to transparency.
At every stage, M2EX provides reasoning to justify its actions, from choosing SQL for retrieval to selecting a bar chart for visualization. Intermediate outputs, like the dataset of war paintings and the Python plotting code, are made visible, ensuring the user can trace the steps taken. The decision making phase wraps up the process by summarizing findings, clarifying the approach, and sharing the final visual result. This shows that M2EX not only answers the query effectively but also ensures its steps are understandable, logical, and well-documented, building trust in its analysis.
Figure 4: Optimization of M2EX: Smart replanning.
Figure 5: Optimization of M2EX: Parallel planning.
Example 2 - Smart Replanning: What is depicted on the oldest Renaissance painting in the database? (see Figure 4).
Contrary to the previous example, M2EX here involves smart replanning - a major optimization technique of M2EX. The main idea is to dynamically adapt the planning in case some tasks of the workflow fail or do not produce any results. Here’s a breakdown of each step:
1) Planning & Expert Model Allocation: M2EX outputs the initial workflow plan that has 2 tasks. The first task involves retrieving the image path and the year of the oldest Renaissance painting in the database using a "text2SQL" expert model. It also involves an "image_analysis" expert model in the second task, which aims to determine what is depicted in the image.
2) Execution and Self-Debugging: M2EX takes the information about the planned workflow as well as task dependencies and puts it into action. In Task 1, it comes with a reasoning statement to generate the SQL query as: SELECT img_path, strftime(’%Y’, inception) AS year FROM paintings WHERE movement = ’Renaissance’ ORDER BY inception ASC LIMIT 1. Then it executes the query over the Artwork database and retrieves the specific image path and year for the oldest Renaissance painting as [’img_path’: ’images/img_0.jpg’, ’year’: ’1438’]. This allows the model to access the actual painting data in the subsequent task.
In Task 2, M2EX utilizes the "image_analysis" expert model (i.e. visual question answering based on BLIP) to examine the contents of img_0.jpg to answer the question: What is depicted in the image? The output of this task is transferred as a final result to the decision making component. At this point, the model’s "thought" process in this component becomes evident. It reasons that while it knows that img_0.jpg is a painting, the details about what is depicted in the painting have not been provided. Therefore, the model decides to not provide a final answer to the user and does replanning.
The replanning capability is a crucial aspect of the M2EX’s approach. Rather than blindly accepting the final answer which does not produce a satisfiable or correct result, the model recognizes the need to replan and calls the "image_analysis" module again. Since the model already knows which image in the database contains the oldest Renaissance painting, it smartly plans the "image_analysis" task as Task 3, by reformulating the question as What is specifically depicted in the painting? M2EX then executes the task, and receives the more concrete answer "umbrellas".
Moving forward, the decision making component confirms the details about the painting. Here, it verifies that the information it has gathered so far aligns with the natural language question and makes sense as a comprehensive understanding of the oldest Renaissance painting. The key aspect is the model’s ability to replan effectively and to strategically leverage the available information to avoid repeating tasks. Example 3 - Parallel Planning: In the Renaissance, find the total number of paintings depicting war and the number of paintings depicting swords (see Figure 5). The figure illustrates how M2EX processes a complex query about Renaissance paintings, focusing on identifying how many paintings depict war and how many depict swords. The pipeline is structured to combine parallel task execution with step-by-step explanations, ensuring clarity and efficiency throughout the process.
The process begins in the Planning & Expert Model Allocation, where the model breaks down the user’s query into distinct subtasks. These subtasks are assigned to specialized modules: Task 1 "text2SQL": This task retrieves image paths and relevant metadata for Renaissance paintings from a database using a SQL query. Task 2 "image_analysis": This task examines whether each painting depicts war. Task 3 "image_analysis": Simultaneously, another module analyzes whether each painting depicts a sword. Task 4 "data_preparation": This task consolidates the results from Task 2 and Task 3 to count and summarize the paintings.
At this point, the model showcases its parallel planning capability. Tasks 2 and 3 are performed concurrently: For Task 2, the system uses image analysis to determine if each painting depicts war. For Task 3, a similar image analysis process identifies paintings that depict swords. Running these tasks in parallel significantly speeds up the workflow, as they operate independently of each other. Once the image analysis tasks are complete, the model transitions to Task 4, where it aggregates the results. The reasoning here details how the system compiles two lists - one for paintings depicting war and one for those depicting swords. Afterwards, M2EX counts the entries in each list. The final results are prepared for the decision making module.
In the decision making phase, the model reflects on its findings. It confirms that sufficient data was processed to answer the query and provides a summary: "There is 1 painting depicting war and 38 paintings depicting swords."
M2EX offers details, explaining how the analysis was conducted and highlighting the disparity between the two categories of paintings. The system further provides an explanation of its methodology, emphasizing how it worked systematically to answer the query. This demonstrates M2EX’s ability to manage tasks efficiently through parallel execution and to ensure transparency through reasoned explanations at every step. By combining these capabilities, the system provides a clear, accurate, and well-supported response to the user’s query.
Note that we did not compare M2EX with NeuralSQL on ArtWork dataset, as such a comparison would be unfair due to NeuralSQL’s inability to support plotting.
Appendix E Error Analysis
(a) CAESURA (ArtWork)
(b) M2EX (ArtWork)
(c) CAESURA (RotoWire)
(d) M2EX (RotoWire)
(e) M2EX (EHRXQA)
@@ Appendix E Error Analysis
Figure 6: Error analysis on different datasets: (a) CAESURA on ArtWork, (b) M2EX on ArtWork, (c) CAESURA on RotoWire, (d) M2EX on RotoWire, and (e) M2EX on EHRXQA.
Error Analysis on the ArtWork Dataset
@@
For 11 errors in multi-modal questions, five are related to single-value outputs, four to plots, and three to data structures. All of these errors are attributed to incorrect outputs generated by the image analysis model. After further research, we found two ambiguous tasks in classifying the error categories. (1) Plot the number of paintings that depict war for each year and (2) What is depicted on the oldest religious artwork in the database? Both tasks failed due to improperly parsed sub question for the image analysis task, specifically the oversimplified term “war.” While this term is semantically related to the correct natural language question, “Does the image depict war?”, it does not fully capture the intent of the task. As a result, it cannot be classified as a completely faulty question. Notably, the M2EX model generated correct results for these tasks, underscoring the limitations of CAESURA’s approach in handling subtle semantic distinctions.
@@
In questions which require a parallel workflow - including two data structures, plot | plot, and plot | data structure outputs — errors are observed at the early planning stage. Our analysis reveals that CAESURA encounters significant challenges in generating accurate plans for embarrassingly parallel tasks. For two of these tasks, the system fails to generate any plan at all. For the remaining four tasks, CAESURA can provide partial results for some subtasks, but other subtasks are left unanswered, reflecting a broader issue in its ability to manage parallel planning. Our M2EX system successfully generates the appropriate plans for all tasks, as shown in Figure 6 (b). In addition, all text-to-SQL steps, data preparation pipelines, and plot outputs, where required, are validated as correct. As illustrated in Figure 6(b), the only source of errors is the inaccurate output of the image analysis model, which accounted for 11 errors. No other errors are located in the text-to-SQL task, plot generation, or task planning deficiencies. This analysis highlights the image analysis model as the bottleneck in system performance, underscoring the need for further refinement in its predictive accuracy.
Error Analysis on the RotoWire Dataset
Figure 6 (c) reveals that CAESURA encounters 9 errors across 12 inference tasks on the RotoWire dataset. These tasks are evenly divided between single-modal and multi-modal categories. Among the three single-modal tasks, one stumbles due to an SQL query missing essential filter clauses, resulting in inaccurate structured data. The other two, focused on plotting, fail to generate visualizations consistent with the analytical findings.
In the multi-modal group, six tasks face challenges. A task requiring a single-value output is derailed by suboptimal text analysis. Additionally, the Bart model’s limited text comprehension hampers two tasks expecting data structure outputs and two others involving plots, all undermined by faulty text interpretation. Another task, aimed at producing a structured output, falters during the planning stage because the strategy cannot be refined within the permitted attempts.
In contrast, our M2EX system, as illustrated in Figure 6 (d), excels by devising suitable plans for all tasks and accurately resolving every single-modal task. However, it encounters issues in four multi-modal tasks: two demanding data structures and one plotting task succumb to flawed text analysis, while a fourth task needing a structured output fails during post-data preparation. Beyond these, no errors arise in text-to-SQL conversions or plot generation. This comparison underscores M2EX’s greater resilience while highlighting text analysis as a shared weakness. CAESURA, however, suffers from additional pipeline limitations.
Error Analysis on the EHRXQA Dataset
@@
Since NeuralSQL is a one-step approach lacking task planning and explainability, we are unable to localize the source of errors as systematically as in the M2EX or CAESURA systems. Consequently, we focus our error analysis solely on the M2EX system using the EHRXQA dataset.
@@
Figure 6 (e) presents the distribution of 49 errors across various steps, categorized by their respective scopes: Image Single-1 (23 errors), Image Single-2 (17 errors), and Image+Table Single (9 errors). Among these, 36 errors are associated with the categorical scope, with 20 attributed to Image Single-1 and 16 to Image Single-2. In contrast, errors linked to the binary output type are primarily found in the Image+Table Single scope. Specifically, Image Single-1 contributes three binary errors, Image Single-2 accounts for one, and Image+Table Single includes nine, summing up to 13 binary errors out of the total 49. Considering the uneven distribution of errors across various output types and scopes, we identified inaccurate image analysis — primarily driven by the M3AE model (Chen et al., 2022) — as the main source of errors. Our analysis reveals that errors linked to categorical output types (36) are nearly three times higher than those associated with binary output types (13). This suggests that the error pattern is less related to the task difficulty across different scopes and more influenced by the output type, as binary questions demonstrate a statistically higher success rate compared to categorical ones. Notably, the Image + Table Single scope exclusively utilizes binary output types.
@@
To gain a deeper understanding, a step-by-step error analysis reveals that out of the 23 errors in the Image Single-1 scope, 22 are due to inaccuracies in image analysis, while only one is related to a misstep in the text-to-SQL process. The specific question text for this case is: “Catalog all the anatomical findings seen in the image, given the first study of patient 11801290 on the first hospital visit.” The generated SQL query fails to include the condition specifying the first study, resulting in an incorrect output. In the Image Single-2 category, 16 out of 17 total errors are due to inaccurate image analysis, with one error attributed to the text-to-SQL step. The specific query in question is: “Does the second-to-last study of patient 16345504 this year reveal still-present fluid overload/heart failure in the right lung compared to the first study this year?”. The text-to-SQL task fails to correctly retrieve the first and last study of this year as required, instead erroneously returning multiple studies from the current year. In the Image+Table Single scope, all nine errors involve binary output types. Of these, six result from inaccurate image analysis, one from incomplete planning, and two from an incorrect text-to-SQL step. The error caused by incomplete planning occurs with the question: “Did patient 19055351 undergo the combined right and left heart cardiac catheterization procedure within the same month after a chest x-ray revealed any anatomical findings until 2104?”. In this case, the plan omits the necessary image analysis step, leading to an incorrect final output. During the reasoning stage, instances were identified where an empty output produced a no response that coincidentally aligned with the ground truth. However, M2EX’s explainability highlights this as a misclassification, as the absence of output was not due to correct reasoning.
@@
Two errors in the Image+Table Single category are attributed to text-to-SQL misbehavior. The specific questions causing these errors are: "Was patient 12724975 diagnosed with hypoxemia until 1 year ago, and did a chest x-ray reveal any tubes/lines in the abdomen during the same period?” and "Was patient 10762986 diagnosed with a personal history of tobacco use within the same month after a chest x-ray showing any abnormalities in the aortic arch until 1 year ago?" In both cases, the SQL queries fail to correctly apply the condition (since current time) until 1 year ago, instead treating 1 year ago as a fixed point in time.